ELECTRONIC TRANSITIONS OF RUTHENIUM MONOXIDE

NA WANG, \underline{Y. W. NG} , and A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

The electronic transition spectrum of ruthenium monoxide (RuO) molecule in the spectral region between 545nm to 640nm has been recorded and analyzed using laser ablation/reaction free-jet expansion and laser induced fluorescence spectroscopy. The RuO molecule was produced by reacting laser- ablated ruthenium atoms with N_{2}O seeded in argon. Nine vibrational bands were recorded and they are identified to be belonging to four electronic transition systems, namely the [18.1]Ω = 4 - X^{5}Δ_{4} transition, [16.0]^{5}Φ_{5} - X^{5}Δ_{4} transition, [18.1]Ω = 3 - X^{5}Δ_{3}, and [15.8] ^{5}Φ_{4} - X^{5}Δ_{3} transition. RuO has been determined to have a X^{5}Δ_{4} ground state. A least squares fit of the measured rotational lines yielded molecular constants for the ground and the low-lying electronic states. A molecular orbital energy level diagram has been used to help with the assignment of the observed electronic states.